OpenStudy (gabylovesyou):

factor x2 − 16x + 64 is it (x - 8) (x - 8)

jimthompson5910 (jim_thompson5910):

right again, you're on a roll

jimthompson5910 (jim_thompson5910):

-8 + (-8) = -16 -8 times -8 = 64 that's a way to check it